OpenStudy (anonymous):

what voltage (V1) is required to produce a total current of 7.5mA.... R1=2.0 kohms, R2=2.0 kohms, R3=2.0 kohms.. (Parallel circuit)

OpenStudy (anonymous):

\[Voltage = I \times R total\] \[1/R = 1/R1 + 1/R2 +1/R3\]

OpenStudy (anonymous):

hey moemoe

OpenStudy (anonymous):

HI

OpenStudy (anonymous):

can u do that , i assume you can

OpenStudy (anonymous):

1.5

OpenStudy (anonymous):

5.025 mA